3/x-6+7/x-2=10/x-4 solve it
ans is 9

Answer by Boreal(15235)

3/(x-6)+7/(x-2)=10/(x-4); multiply everything by (x-6)(x-2)(x-4)
3(x-2)(x-4)+7(x-4)(x-6)=10(x-2)(x-6)
3(x^2-6x+8)+7(x^2-10x+24)=10(x^2-8x+12)
3x^2-18x+24+7x^2-70x+168)=10x^2-80x+120
10x^2-88x+192=10x^2-80x+120
-8x=-72
x=9
check 1+1=2